A number, y, is equal to twice the sum of a smaller number and 3. The larger number is also equal to 5 more than 3 times the smaller number. Would it be 2x-y=6 and 3x-y=-5
Answer by macston(5194)

.
y=twice the sum of smaller number plus 3
y=2(x+3)
y=2x+6
y=5 more than 3 times x
y=3x+5
2x+6=3x+5
1=x
.
y=2(x+3)
y=2(1+3)
y=2(4)=8
.
CHECK:
y=3x+5
8=3(1)+5
8=3+5
8=8